Consider a coaxial cable formed by two concentric cylindrical conductors with the radius of the inner conductor being “a” and the outer conductor being “b”. Let L be the length of the cable. Assume the surface charge distribution on the inner conductor be defined by ρs. Determine the flux density D at a point ρ located radially between a and b, i.e. a < ρ < b. 


1
Expert's answer
2020-09-24T10:24:01-0400

"ds = ad\phi dz"

"Q=\iint Psad\phi dz"

"= 2\pi PsL" ---------------- 1

from gauss law,

"Q= DsA"

Change Q at point P located is

"Q = Ds \times 2\pi PL" ------------ 2

from equation 1 and 2

"Q= 2\pi aPsL"

"2\pi PsL =" "Ds \times 2\pi PL"

"Ds = a\frac{Ps}{p}"

(a<P<b)